# Text2Cypher Across Languages: Evaluating Foundational Models Beyond English

### v1

@@ 4.1. Impact of question language
Tarea: Generar una sentencia Cypher para consultar una base de datos de grafos. Instrucciones: Usa únicamente los tipos de relaciones y propiedades proporcionados en el esquema. No utilices ningún otro tipo de relación ni propiedades que no estén incluidas en el esquema. No incluyas explicaciones ni disculpas en tus respuestas. No respondas a ninguna pregunta que no sea una solicitud para construir una sentencia Cypher. No incluyas ningún texto excepto la sentencia Cypher generada. Görev: Bir çizge veritabanını sorgulamak için bir Cypher ifadesi oluştur. Talimatlar: Yalnızca şemada verilen ilişki türlerini ve özellikleri kullan. Şemada verilmeyen herhangi bir ilişki türünü veya özelliği kullanma. Yanıtlarında hiçbir açıklama veya özür ifadesine yer verme. Cypher ifadesi oluşturmak dışında başka bir şey isteyen sorulara yanıt verme. Oluşturulan Cypher ifadesi dışında hiçbir metin ekleme.
@@
4.2. Impact of prompt language
@@
We also translated the prompts into Spanish and Turkish to evaluate the impact of prompt language on model performance. The prompts were translated using the GPT-4o-mini model and are shown in Table 3. The performance comparison for Spanish and Turkish with and without translated prompts are presented in Figure 3. The results show that when prompts are translated, the ROUGE-L scores, which used for translation-based evaluation comparing ground-truth and generated text, improve slightly (by about 1–1.5%). However, Exact-Match scores for execution-based evaluation remain largely unchanged. These results suggest that prompt translation has a minor impact on the overall result in our setup. It is important to note that in our dataset, the database schemas, including nodes, relationships and properties, remain in English. Future work could explore the impact of localizing schema items to match the language of the input and prompt, which may reveal further interactions between language and performance.
(a) Translation-based evaluation: ROUGE-L score
(b) Execution-based evaluation: Exact-Match score
Figure 3. Model performance with/out prompt translation.
5. Conclusion
@@
This work presents a performance comparison of foundational LLMs on the Text2Cypher task across English, Spanish, and Turkish. We created and released a multilingual test set by translating English questions into Spanish and Turkish while keeping the same ground-truth queries, enabling fair cross-lingual comparison of model performance. Our results consistently show that foundational models perform best on English questions, followed by Spanish, and lowest on Turkish. This pattern is likely driven by differences in training data availability and the linguistic characteristics, which pose additional challenges for the models. We also explored the impact of translating task instruction prompts alongside the input questions. Our findings suggest that prompt translation slightly improves translation-based evaluation but has little to no effect on query execution results, suggesting that prompt language has minimal impact on overall performance in our setting. Future work includes schema localization to improve multilingual performance, fine-tuning models on individual or combined languages and expanding evaluation to additional languages.

### v2

@@ 4.2 Impact of question language
Tarea: Generar una sentencia Cypher para consultar una base de datos de grafos. Instrucciones: Usa únicamente los tipos de relaciones y propiedades proporcionados en el esquema. No utilices ningún otro tipo de relación ni propiedades que no estén incluidas en el esquema. No incluyas explicaciones ni disculpas en tus respuestas. No respondas a ninguna pregunta que no sea una solicitud para construir una sentencia Cypher. No incluyas ningún texto excepto la sentencia Cypher generada.
@@
TR System
Instruct.
Görev: Bir çizge veritabanını sorgulamak için bir Cypher ifadesi oluştur. Talimatlar: Yalnızca şemada verilen ilişki türlerini ve özellikleri kullan. Şemada verilmeyen herhangi bir ilişki türünü veya özelliği kullanma. Yanıtlarında hiçbir açıklama veya özür ifadesine yer verme. Cypher ifadesi oluşturmak dışında başka bir şey isteyen sorulara yanıt verme. Oluşturulan Cypher ifadesi dışında hiçbir metin ekleme.
User
Instruct.
@@
(a) Translation-based: ROUGE-L score
(b) Execution-based: Exact-Match score
Figure 3: Model performance with/out prompt translation.
4.3 Impact of prompt language
@@
We also translated the prompts into Spanish and Turkish to evaluate the impact of prompt language on model performance. The prompts were translated using the GPT-4o-mini model and are shown in Table 3. The performance comparison for Spanish and Turkish with and without translated prompts are presented in Figure 3. The results show that when prompts are translated, the ROUGE-L scores, which used for translation-based evaluation comparing ground-truth and generated text, improve slightly (by about 1–1.5%). However, Exact-Match scores for execution-based evaluation remain largely unchanged. These results suggest that prompt translation has a minor impact on the overall result in our setup.
It is important to note that in our dataset, the database schemas, including nodes, relationships and properties, remain in English. Future work could explore the impact of localizing schema items to match the language of the input and prompt.
4.4 Impact of finetuning
After observing performance differences across languages, we next explore how finetuning affects results. Among the evaluated models, we selected Meta-Llama-3.1-8B-Instruct for fine-tuning due to its multilingual support, open availability, and compatibility with our Unsloth-based setup. Considering the time and resource constraints, we focused on a single model and considered fine-tuning others as a future work. We finetune the selected baseline model separately on two datasets:
  • English-only finetuned model: Trained on the English-only dataset described in Section 3.3. This setup allows us to evaluate how English-only finetuning affects performance on both English and non-English inputs.
  • Multilingual finetuned model: Trained on the multilingual dataset to assess the benefits of exposure to multiple languages during finetuning.
(a) Translation-based: ROUGE-L score
(b) Execution-based: Exact-Match score
Figure 4: Comparison of model performance: Foundational model (blue), Fine-tuned model on English-only data (orange), and Fine-tuned model on multilingual data (green).
The finetuned models are evaluated on the test set using data per language. In Figure 4 we present the performance of the foundational model, finetuned model on English-only dataset and finetuned model on multilingual dataset. The figure shows that using the English-only dataset improves model performance of the Text2Cypher task across all languages. However, the improvement for Turkish is smaller compared to English and Spanish. For example, based on the translation-based ROUGE-L score, English and Spanish improve by around 0.20, while Turkish improves by approximately 0.15. Spanish still performs worse than English, followed by Turkish. This performance gap may be attributed to linguistic differences: English and Spanish belong to the Indo-European family, whereas Turkish is an Altaic language with different characteristics. When evaluating the model finetuned on the multilingual dataset, we observe that performance across languages becomes more balanced. For instance, ROUGE-L scores reach 0.86 for English, 0.85 for Spanish, and 0.83 for Turkish. These results suggest that multilingual finetuning improves generalization and narrows the performance gap across languages.
Overall, these results indicate that both language resourcedness and linguistic characteristics influence model performance. Foundational models and those finetuned only on English struggle with non-English inputs, particularly for languages with different characteristics or with less resources. Multilingual finetuning mitigates these challenges and leads to more balanced performance across languages.
5 Conclusion
@@
This work presents a performance comparison of foundational and finetuned LLMs on the Text2Cypher task across English, Spanish, and Turkish. We created and released a multilingual dataset by translating user questions in English into Spanish and Turkish, enabling fair cross-lingual evaluation. Our results consistently show that foundational models perform best on English inputs, followed by Spanish, and then Turkish. This pattern is likely driven by differences in language resourcedness and linguistic characteristics, which introduce additional challenges for LLMs. We also explored the effect of translating task instruction prompts alongside the input questions. Our findings suggest prompt language has minimal impact on overall performance in our setting. Finally, we finetuned a foundational model on both English-only and multilingual datasets. The results revealed that English-only finetuning improves performance across all languages, it does not eliminate the performance gap. In contrast, multilingual finetuning improves generalization and narrows the performance gap across languages. Future directions include schema localization to further support multilingual performance, expanding evaluation to additional languages, and exploring alternative alignment strategies beyond finetuning to improve cross-lingual robustness.
Declaration on Generative AI Usage
During the preparation of this work, the author(s) used Chat-GPT in order to: ’Improve writing style’ and ’Paraphrase and reword’. After using these tool(s)/service(s), the author(s) reviewed and edited the content as needed and take(s) full responsibility for the publication’s content.